### v1

@@ Discussion
In conclusion, our framework represents a significant advancement in the field of XAI, offering a systematic and generic approach to optimizing the explainability of deep learning models. By providing optimally faithful and comprehensible explanations, our framework enhances trust and interpretability, paving the way for broader adoption of AI technologies in critical domains such as healthcare and medicine.
Methods
Evaluation metrics for the explanation
A crucial aspect of this study lies in evaluating ”how accurate and comprehensive is an explanation?” To derive a useful explanation, two primary scores play a pivotal role: fidelity and complexity. An intuitive way to assess the quality of an explanation is by measuring its ability to accurately capture how the predictive model reacts to random perturbations [34]. For a deep neural network , and input features , the feature importance scores (also known as ”attribution scores”) are derived in a way such that when we set particular input features to a baseline value , the change in the network’s output should be proportional to the sum of attribution scores of the perturbed features . We quantify this by using the Pearson correlation between the sum of the attributions of and the difference in the output when setting those features to a reference baseline [35]. Thus, we define the faithfulness of an explanation method as:
where is a subset of indices (), is a sub-vector of an input ( and the unchanged features of image). The total number of the sub-vectors, which partition an image is . We denote as the changed features, and the unchanged features of image.
If for an image , explanation highlights all features, then it may be less comprehensible and more complex than needed (especially for large ). It is important to compute the level of complexity, as an efficient explanation has to be maximally comprehensible [35]. If is a valid probability distribution and the is the fractional contribution of feature to the total magnitude of the attribution, then we define the complexity of the explanation for the network as:
where:
In order to evaluate these two explainability metrics, we used the software developed by [24]. This software package is a comprehensive toolkit that collects, organizes, and evaluates a wide range of performance metrics, proposed for explanation methods. We note that we used a zero baseline (’black’; = ), and 70 random perturbations to calculate the Faithfulness score.
Overview of the proposed framework
The proposed framework entails an optimization approach aimed at attaining the highest faithfulness and lowest complexity of an explanation tailored to a specific deep learning network and computer vision task. Key components of the framework include the utilization of baseline XAI methods, a ’Weighted Average’ explanation, a non-linear network (’Non-linear structure’), and an Up-Sampling unit (Fig. 2). The framework incorporates a range of explanations obtained from established XAI methods and computes a adjusted baseline explanation by averaging these explanations based on assigned weights (’Weighted Average’). The assigned weights are computed from the equation below:
where and represent fixed values ranging from 0 to 1, signifying the contributions of faithfulness and complexity, respectively, to the final weighted average. In this study, we analyzed predictions and explanations for ten randomly selected prediction instances from the validation cohort. We assessed the faithfulness and complexity scores for each of the explanation methods and for each prediction instance. Subsequently, we calculated the average faithfulness and complexity scores across the ten prediction instances and determined the final assigned weights using eq. 4 with . The average values of , and the assign weights were normalized to range from 0 to 1.
The explanations generated by the baseline XAI methods are fed into a non-linear network consisting of an encoder, bottleneck, and decoder architecture (see later sections for more details), with the objective of reconstructing a unique explanation. Following this, the reconstructed explanation undergoes sequential processing through an up-sampling deconvolution layer to double resolution. In our case, the resolutions of the 3D inputs in the application were , while for the 2D images, they were . The final output of our explanation of optimizing framework consists an up-sampled higher-resolution explanation based on the input ( and ) and an explanation with the same resolution as the input (, and ). The up-sampling approach is motivated by the aim to enhance the detail of the explanation. We note that particularly in neuroscience and medical applications, a higher resolution is closely linked with increased trust of the community to AI procedures, making the provision of explanations in both higher and intial resolutions highly desirable.
where represents the derived explanation by our optimizer, denotes the ’Weighted Average’ explanation, indicates the average of , signifies the variance of , represents the covariance of and , and and are two variables utilized to stabilize the division with a weak denominator [36]. refers to the similarity score between the ’Weighted Average’ explanation and the down-sampled predicted explanation (low resolution), while signifies the similarity score between a tri-linear or bicubic (in 3D and 2D space respectively) up-sampled ’Weighted Average’ explanation and the derived high-resolution explanation. The values of and were fixed at 0.5 each. The total loss function was given by:
Following an ablation study exploring various combinations of ’s in the above cost function, we determine that setting and is optimal. This selection ensures a balanced optimization approach that effectively addresses the unique challenge of deriving an optimal explanation.
The experiments in 3D and 2D applications
In the 3D application, our focus is on a binary classification task concerning the presence or absence of the paracingulate sulcus (PCS). Leveraging a comprehensive cohort of 596 subjects from the TOP-OSLO study [25], we divide the dataset into training (60%), validation (20%), and testing (20%) sets. For the 3D classification task, we utilize white-grey surface matter images extracted from T1-weighted MRI scans of patient brains. Our image preprocessing follows the outlined steps in [5]. For the 2D application, we utilize the publicly available CIFAR-10 dataset (https://paperswithcode.com/dataset/cifar-10). This dataset, a subset of the Tiny Images dataset, consists of 60,000 color images with dimensions of . Each image is associated with one of 10 mutually exclusive classes, such as airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. The dataset is organized into training (60%), validation (20%), and testing (20%) sets, with 6,000 images per class.
Deep learning architectures
In this study, we utilized a Resnet-18 and a ResNet-50 network architectures [37] for the classification tasks in three-dimensional (3D) and two-dimensional (2D) settings, respectively.
With regards to the explanation optimizer, for the 3D application, we employed the Swin-UNETR architecture [38] as our non-linear reconstruction network, featuring a 24-feature size (Fig. 2 a.). For the 2D explanation optimization, we utilized a U-NET encoder-decoder architecture [39] (Fig. 2 b.).
As part of an ablation analysis (see Results), we explored a linear implementation in the three-dimensional application. Here, we employed visual transformer encoders [40] to estimate weights for each of the explainable baseline methods (refer in Supplementary material subsection: ’Deep learning architectures: The linear implementation of the proposed network’). Subsequently, these estimated weights were utilized in a weighted average to extract the linearly optimized explanation.
The baseline XAI methods
For this study, we employed eight established XAI methods: Saliency[20], DeepLift[20], Kernel Shap[20], DeepLift Shap[20], Integrated Gradients [41], Guided Backpropagation [42], Guided GradCam [43] and Gradient Shap. GradientShap makes an assumption that the input features are independent and that the explanation model is linear, meaning that the explanations are modeled through the additive composition of feature effects. Under these assumptions, SHAP values [20] can be approximated as the expectation of gradients computed for randomly generated samples from the input dataset. Gaussian noise is added times to each input, creating different baselines for computing the SHAP values.
Extending the implementation of these explainability approaches from two-dimensional (2D) to three-dimensional (3D) space was necessary. We utilized the Python software developed by [24] as the underlying libraries for our implementation. Further implementation details can be found in the public GitHub repository associated with this study. It’s important to note that all explanations were modified to focus exclusively on positive attributions. Any negative attribution in the original explanations were replaced with a zero value.
The hyperparameter utilization
For both the 2D and 3D classification tasks, after randomly shuffling the data, each dataset underwent splitting into training, validation, and testing sets, with proportions of , , and of the total number of images, respectively. Sparse categorical cross-entropy was employed as the cost function, while the Adam algorithm [44] was utilized to optimize the loss function. Initially, the learning rate remained fixed for the first 300 epochs, followed by a decrease of every 100 epochs. To prevent overfitting, an early stopping criterion of 100 consecutive epochs was enforced, with a maximum of 500 epochs used for the input modalities in the corresponding 2D and 3D applications. Data augmentation techniques were applied, including rotation (around the center of the image by a random angle from the range of ), width shift (up to pixels), height shift (up to pixels), and Zero phase Component Analysis (ZCA;[45]) whitening (adding noise to each image). Hyperparameter tuning was conducted for both applications using four different values of initial learning rates: , , , and for the 2D application, and , , , and for the 3D application.
The same fixed step learning rate and optimization algorithm, namely the Adam algorithm [44], were employed to optimize the loss function in the explainability task. However, no data augmentation techniques were utilized to train the optimizer. The cost function used was that of eq. 7. For the 2D task, the maximum number of epochs was set at 150, with an early stopping criterion of 10 epochs applied after passing the first 90 epochs. For the 3D task, the maximum number of epochs was set at 500, with an early stopping criterion of 120 epochs applied after the initial 300 epochs. Once again hyperparameter tuning was conducted for both applications using four different values of initial learning rates: , , , and for the 2D application, and , , , and for the 3D application.
Code and dataset availability
@@ Methods
Acknowledgment All research at the Department of Psychiatry in the University of Cambridge is supported by the NIHR Cambridge Biomedical Research Centre (NIHR203312) and the NIHR Applied Research Collaboration East of England. The views expressed are those of the author(s) and not necessarily those of the NIHR or the Department of Health and Social Care. We acknowledge the use of the facilities of the Research Computing Services (RCS) of University of Cambridge, UK. GKM consults for ieso digital health. All other authors declare that they have no competing interests.
Author information
Authors and Affiliations
Department of Psychiatry, University of Cambridge, Cambridge, UK.
Michail Mamalakis, Graham Murray, John Suckling
Department of Enviromental Science, University of Virginia, Charlottesville, Virginia, United States.
Antonios Mamalakis
Department of Computer Science and Technology, Computer Laboratory, University of Cambridge, Cambridge, UK.
Michail Mamalakis, Pietro Lio
Department of Psychiatric Research, Diakonhjemmet Hospital, Oslo, Norway
Ingrid Agartz
Norment, Division of Mental Health and Addiction, Oslo University Hospital, Institute of Clinical Medicine, University of Oslo, Oslo, Norway.
Lynn Egeland Mørch-Johnsen
Department of Psychiatry and Department of Clinical Research, Østfold Hospital, Grålum, Norway.
Lynn Egeland Mørch-Johnsen
Contributions M.M, A.M., P.L. conceived the study. M.M. wrote the code and conducted the experiments. I.A., LE.MJ. and M.M. collected and pre-processed the data cohort. M.M, and A.M. analyzed the data and results. M.M. contributed to pulling deep learning and XAI methods and conducted chart reviews. M.M contributed to the experimental design and validation protocol. A.M. M.M. and P.L. was in charge of overall direction and planning. All authors contributed to the interpretation of the results. M.M., and A.M. visualized the study and extracted figures, and M.M and A.M. drafted the manuscript, which was reviewed, revised and approved by all authors.
Corresponding author Correspondence to Michail Mamalakis or Pietro Lio
Ethics declarations
Competing interests
The authors declare no competing interests.

### v2

@@ 5 Discussion
6 Conclusions
The sub-optimal results of individual XAI methods, combined with high inter-method variability, present a significant challenge in the field of XAI, as this variability undermines trust in deep network predictions. This study addresses this challenge by introducing a novel framework designed to enhance explainability through the integration of multiple established XAI methods. Central to our approach is a non-linear neural network model, the ’explanation optimizer,’ which synthesizes a singular, optimal explanation from these methods. Crucially, this process goes beyond a simple combination or ensembling of existing XAI tools. Instead, it employs a carefully designed nonlinear, constrained optimization technique, leveraging a non-trivial loss function to achieve results that outperform those produced by the individual methods or their straightforward combinations.
The optimizer operates with two key metrics that evaluate the quality of the generated explanations: faithfulness, which measures how accurately the explanation mirrors the network’s decision-making, and complexity, which assesses the comprehensibility of the explanation for end users. By carefully balancing these metrics, the optimizer generates explanations that are both accurate and accessible. Our experimental validation across multi-class and binary classification tasks in both 2D object and 3D neuroscience imaging demonstrates the efficacy of this approach. Notably, our explanation optimizer achieved superior faithfulness scores, with an average increase of 155% and 63% over the best-performing individual XAI methods in 3D and 2D applications, respectively, while also reducing complexity to enhance comprehensibility.
This framework represents a significant advancement in XAI by systematically enhancing the explainability of deep learning models. By generating explanations that are both highly faithful and easy to interpret, our approach addresses the long-standing issue of inter-method variability, strengthens trust and interpretability, and promotes the broader adoption of AI technologies in critical fields such as healthcare and medicine.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment